OpenStudy (anonymous):

2 ≤ | x - 2 | inequalty

OpenStudy (babalooo):

yup, that sure is inequality equation right there

OpenStudy (anonymous):

\(x\ge4\) or \(x\le0\) to satisfy the inequality

OpenStudy (texaschic101):

2 <= | x + 2 | -2 <= x + 2 <= 2 --- subtract 2 -2 - 2 <= x <= 2 - 2 -- combine like terms -4 <= x <= 0 so we have : x <= -4 and x <= 0

OpenStudy (alekos):

I get x>=4 and x<=0 same as Orion

OpenStudy (babalooo):

2<= |x-2| 2<=x-2 4<=x 2 <= -x+2 0<= -x 0 >=x

OpenStudy (alekos):

yep. that's it